OpenStudy (anonymous):

Solve the following equation: x2 - 14x + 49 = 0 {7} {-7} {-7, 7} { 7,-7}

OpenStudy (anonymous):

huh. (x-7)^2 = 0 so x = 7. equal roots.

OpenStudy (anonymous):

x2 - 14x + 49 = 0 x2 - 7x-7x + 49 = 0 (x-7)^2=0 x=7

OpenStudy (anonymous):

medal plzz:):):)

OpenStudy (anonymous):

can i also have the medal???